# Total variation distance between SDEs with stable noise and Brownian motion with applications to Poisson PDEs

### v1

@@ Abstract.
We consider a $d$-dimensional stochastic differential equation (SDE) of the form $dU_{t}=b(U_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}Z_{t}$ with initial condition $U_{0}=u\in{\mathds{R}^{d}}$ and additive noise. Denote by $X_{t}^{x}$ the (unique) solution with initial value $X_{0}^{x}=x$, if the driving noise $Z_{t}$ is a $d$-dimensional rotationally symmetric $\alpha$-stable ($21<\alpha<$ 2) Lévy process, and by $Y_{t}^{y}$ the (unique) solution with initial value $Y_{0}^{y}=y$, if the driving noise is a $d$-dimensional Brownian motion.
We give precise estimates of the total variation distance $\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{y})\|_{\mathrm{TV}}$ for all $t>0$, and we show that the ergodic measures $\mu_{\alpha}$ and $\mu_{2}$ of $X_{t}^{x}$ and $Y_{t}^{y}$, respectively, satisfy $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right]$.
As an application, we consider the following $d$-dimensional Poisson PDE
$-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
where $h\in C_{b}({\mathds{R}^{d}},\real)$ and $1<\alpha\leq 2$. We show that this PDE admits a solution $f_{\alpha}$, which is continuous, but may have linear growth. For fixed $\alpha_{0}\in(1,2)$, we show that for any $\alpha_{0}\leq\alpha<2$ there is a constant $C(\alpha_{0})$ such that
$\left|f_{\alpha}(x)-f_{2}(x)\right|\leq C(\alpha_{0})(1+|x|)\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha}\quad\text{for all $x$}.$
@@ 1. Introduction
This paper can be seen as a continuation of our previous paper [5], where we studied the 1-Wasserstein distance between stochastic differential equations (SDEs) with different drivers based on the methods developed in [1]. This means that we shall work in the same set-up and under similar assumptions, which we will briefly recall.
1.1. General setup and notation
Consider the SDEs
@@ 1.1. General setup and notation
driven by a standard $d$-dimensioal Brownian motion $B_{t}$ and a stable Lévy process $L_{t}$ with characteristic function $\mathds{E}\,\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, $21<\alpha<$ 2; this is a standard rotationally symmetric $\alpha$-stable process run at half speed. The drift coefficient $b:{\mathds{R}^{d}}\to{\mathds{R}^{d}}$ will be specified further down, $\sigma$ is an invertible $d\times d$ matrix.
Our main assumptions are as follows:
@@
There exist constants $\theta_{0}>0$ and $K\geq 0$ such that
$\langle x-y,b(x)-b(y)\rangle\leq-\theta_{0}|x-y|^{2}+K\quad\text{for all\ \ }x,y\in{\mathds{R}^{d}};$
@@
$b\in C^{2}({\mathds{R}^{d}},{\mathds{R}^{d}})$ and there exist constants $\theta_{1},\theta_{2}\geq 0$ such that for all $x,v,w\in{\mathds{R}^{d}}$,
@@
(1.4) $|\nabla_{w}\nabla_{v}b(x)|$ $\leq\theta_{2}|v||w|.$
If $f:{\mathds{R}^{d}}\to\real$ is smooth enough, the directional derivative in direction $v\in{\mathds{R}^{d}}$ is defined as
@@
and $\nabla_{w}\nabla_{v}f(x):=\nabla_{w}(\nabla_{v}f)(x)=\nabla^{2}f(x)v\cdot w$ for $v,w\in{\mathds{R}^{d}}$. The vector and matrix norms of $\nabla f$ and $\nabla^{2}f$ are given by
$|\nabla f(x)|=\sup_{v\in{\mathds{R}^{d}},|v|=1}\left|\nabla_{v}f(x)\right|\quad\text{and}\quad\|\nabla f(x)\|_{\mathrm{op}}=\sup_{v,w\in{\mathds{R}^{d}},|v|=|w|=1}\left|\nabla_{w}\nabla_{v}f(x)\right|$
[…]
@@
$\nabla^{2}f(x)vu$ $=\nabla_{u}\nabla_{v}f(x)=(\nabla_{u}\nabla_{v}f_{1}(x),\dots,\nabla_{u}\nabla_{v}f_{d}(x))^{\top}.$
The Hilbert–Schmidt inner product of two matrices $A,B\in{}^{d\times d}$ is $\left\langle A,B\right\rangle_{\mathrm{HS}}=\sum_{i,j=1}^{d}A_{ij}B_{ij}$, and the Hilbert-Schmidt norm is $\|A\|_{\mathrm{HS}}=\sqrt{\sum_{i,j=1}^{d}A_{ij}^{2}}$.
Finally, $\mathrm{Lip}(\lambda)$ is the family of Lipschitz functions $f:{\mathds{R}^{d}}\to\real$ with Lipschitz constant $\lambda$ and norm $\|f\|_{\mathrm{Lip}}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|}\leq\lambda$.
We will frequently need the constants
(1.5) $A(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{d/2}\Gamma(1-\frac{\alpha}{2})},\quad\omega_{d-1}:=\frac{2\pi^{d/2}}{\Gamma\left(\frac{d}{2}\right)},$
which can be estimated as follows (see [5, Lemma 3.3]):
Under (1.3) the both SDEs (1.1) and (1.2) have unique non-explosive (strong) solutions, which we denote by $X_{t}$ and $Y_{t}$, respectively. These are time homogeneous Markov processes and write $X_{t}^{x}$ and $Y_{t}^{y}$ to indicate the initial position $X_{0}^{x}=x$ and $Y_{0}^{y}=y$. The operator semigroup induced by the Markov process $(X_{t}^{x})_{t\geq 0}$ is
@@
From the classical Lyapunov function criterion, see e.g. [10], we know that under (H1) and (H2) the solutions to the SDEs (1.1) and (1.2) are ergodic. Denote by $\mu_{\alpha}$ and $\mu_{2}$ the respective ergodic measures of $X_{t}^{x}$ and $Y_{t}^{y}$.
@@ 1.2. Main results
We will now describe the principal results of our paper. Write $\mathscr{C}_{b}(1)=\{h\in C_{b}({\mathds{R}^{d}},\real)\>;\>\|h\|_{\infty}\leq 1\}\}$ for the unit ball in $C_{b}({\mathds{R}^{d}},\real)$. The total variation distance of two probability measures (on ${\mathds{R}^{d}}$) $\mu$ and $\nu$ is given by
(as usual we write $\mu(h):=\int h\,\mathrm{d}\mu$ etc.). It is clear that $\|\mu-\nu\|_{\mathrm{TV}}\leq 2$ for all probability measures $\mu$ and $\nu$.
Throughout this paper, $C,C_{1},C_{2}$ denote positive constants which may depend on $\theta_{0},\theta_{1},\theta_{2},K,\|\sigma\|_{\mathrm{op}},\|\sigma^{-1}\|_{\mathrm{op}}$, but they are always independent of $d$ and $\alpha$; their values may change from line to line.
@@
$\qquad\leq 2\wedge\left[C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|+\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
@@
(1.7) $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\right].$
In the estimate (1.7) we cannot hope, in general, for a faster speed of convergence than $2-\alpha$. This can be seen from the one-dimensional Ornstein–Uhlenbeck, cf. Proposition 7.1 below.
We can apply this theorem to study the convergence rate of Poisson PDEs as follows ($1<\alpha\leq 2$):
(1.8) $-(-\Delta)^{\alpha/2}f(x)+\langle b(x),\nabla f(x)\rangle=h(x)-\mu_{\alpha}(h),$
where $h\in C_{b}({\mathds{R}^{d}},\real)$,
$-(-\Delta)^{\alpha/2}f(x)=\int_{{\mathds{R}^{d}}\setminus\{0\}}\left[f(x+z)-f(x)-\langle\nabla f(x),z\rangle\mathds{1}_{(0,1)}(|z|)\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z$
is the fractional Laplacian for $\alpha\in(1,2)$, and $\Delta f(x)=\sum_{i=1}^{d}\partial^{2}_{i}f(x)$ is the usual Laplace operator. This type of Poisson equation has many applications in central limit theorems related to SDEs and approximations of probability measures, see [9] and [7, 20].
Our second theorem is as follows:
Theorem 1.2.
Assume (H1) and (H2).
  1. a)
    For any $\alpha\in(1,2]$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the equation (1.8) admits a solution $f_{\alpha}\in C_{\mathrm{lin}}({\mathds{R}^{d}},\real)$.
  2. b)
    Let $\alpha_{0}\in(1,2)$ be arbitrary. For any $\alpha\in[\alpha_{0},2)$ and $h\in C_{b}({\mathds{R}^{d}},\real)$, the solution $f_{\alpha}$ from a) satisfies
    $\|f_{\alpha}-f_{2}\|_{\mathrm{lin}}\leq C(\alpha_{0})\|h\|_{\infty}d^{3/2}(2-\alpha)\log\frac{1}{2-\alpha},$
    with a constant $C(\alpha_{0})$ depending on $\alpha_{0},\theta_{0},\theta_{1},\theta_{2},K$.
2. The key gradient estimate for Theorem 1.1
This section is devoted to proving the following gradient estimate, which is crucial for the proof of our main result.
@@
Assume (H2). For any $t\in(0,1]$, $\alpha\in(0,2)$ and $f\in C^{2}_{b}({\mathds{R}^{d}},\real)$,
[…]
@@ 2.1. A random time-change method
Let $\gamma^{\epsilon}$ be the inverse function of $\ell^{\epsilon}$, i.e., $\ell^{\epsilon}\circ\gamma^{\epsilon}_{t}=t$ for all $t\geq\ell_{0}^{\epsilon}$ and $\gamma^{\epsilon}\circ\ell^{\epsilon}_{t}=t$ for all $t\geq 0$. By construction, $t\mapsto\gamma_{t}^{\epsilon}$ is absolutely continuous on $[\ell_{0}^{\epsilon},\infty)$. Define
@@ 2.2. Integration by parts via Malliavin Calculus
$\langle f,g\rangle_{H}:=\int_{\ell^{\epsilon}_{0}}^{t}\langle f(s),g(s)\rangle\,\mathrm{d}s\quad\text{for $f,g\in H$}.$
@@
Lemma 2.2.
Assume (H2), and fix $\ell\in\mathds{S}$ and $\epsilon\in(0,1)$. For any $t>0$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell^{\epsilon}}_{t}\right)\right]\right|\leq C\left(\frac{1}{\sqrt{\ell^{\epsilon}_{t}-\ell^{\epsilon}_{0}}}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell^{\epsilon}}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
@@
For any fixed $t>\ell^{\epsilon}_{0}$ and $1\leq i\leq d$, let
(2.7) $u_{i}(s):=\frac{1}{t-\ell^{\epsilon}_{0}}\,\sigma^{-1}J_{s,t}^{-1}e_{i}.$
@@
$D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}=e_{i},\quad 1\leq i\leq d.$
Combining this with the duality (2.5) between the Malliavin derivative and the divergence operator gives us, for any $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $1\leq i\leq d$,
$\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]$ $=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),e_{i}\big{\rangle}\right]=\mathds{E}^{\mathds{W}}\left[\big{\langle}\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),D_{u_{i}}Z^{x;\ell^{\epsilon}}_{t}\big{\rangle}\right]$
$=\mathds{E}^{\mathds{W}}\left[D_{u_{i}}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]=\mathds{E}^{\mathds{W}}\left[f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\delta\left(u_{i}\right)\right].$
By the Cauchy–Schwarz inequality and [12, Proposition 1.3.1], we obtain
(2.8) $\begin{split}\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|&\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\delta(u_{i})^{2}\right]}\\ &\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\|u_{i}\|_{H}^{2}+\|Du_{i}\|_{H\otimes H}^{2}\right]},\end{split}$
@@
(2.9) $\begin{split}\|u_{i}\|_{H}^{2}&=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\exp\left(-\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)e_{i}\right|^{2}\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\exp\left(2\theta_{1}\int_{s}^{t}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}}{t-\ell^{\epsilon}_{0}}\,\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}.\end{split}$
Similarly, using (1.3) and (2.6) gives for any $v\in H$
(2.10) $\left|D_{v}Z^{x;\ell^{\epsilon}}_{t}\right|\leq\|\sigma\|_{\mathrm{op}}\|v\|_{H}\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{t-\ell^{\epsilon}_{0}}.$
@@
$\begin{split}D_{v}J_{s,t}^{-1}&=\int_{0}^{1}\exp\left(-\tau\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\left(-\int_{s}^{t}\nabla^{2}b\left(Z^{x;\ell^{\epsilon}}_{r}\right)D_{v}Z^{x;\ell^{\epsilon}}_{r}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\mbox{}\times\exp\left(-(1-\tau)\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}\tau,\end{split}$

### v2

@@ Abstract.
We consider a $d$-dimensional stochastic differential equation (SDE) of the form $\mathrm{d}U_{t}=b(U_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}Z_{t}$, let $X_{t}$ be the solution if the driving noise $Z_{t}$ is a $d$-dimensional rotationally symmetric $\alpha$-stable process ($1<\alpha<2$), and let $Y_{t}$ be the solution if the driving noise is a $d$-dimensional Brownian motion.
Continuing the work in [5], we derive an estimate of the total variation distance $\|\mathrm{law}(X_{t})-\mathrm{law}(Y_{t})\|_{\mathrm{TV}}$ for all $t>0$, and we show that the ergodic measures $\mu_{\alpha}$ and $\mu_{2}$ of $X_{t}$ and $Y_{t}$, respectively, satisfy
$\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
We shall show that this bound is optimal with respect to $\alpha$ by an Ornstein–Uhlenbeck SDE. Combining this bound with a recent interpolation result from [8], we can derive a bound in Wasserstein-$p$ distance ($0<p<1$):
$\|\mu_{\alpha}-\mu_{2}\|_{W_{p}}\leq\frac{Cd^{(p+3)/2}\log(1+d)}{\alpha-1}(2-\alpha).$
Key Words: Total variation distance, Wasserstein-$p$ distance, stochastic differential equation, Poisson equation, stable process.
@@ 1. Introduction
Let us consider the SDEs
@@
driven by a standard $d$-dimensioal Brownian motion $B_{t}$ and a stable Lévy process $L_{t}$ with characteristic function $\mathds{E}\,\mathrm{e}^{\mathop{\mathrm{i}}\xi L_{t}}=\mathrm{e}^{-t|\xi|^{\alpha}/2}$, $1<\alpha<2$; this is a standard rotationally symmetric $\alpha$-stable process run at half speed. Our main assumptions are as follows:
@@
$\sigma\in\real^{d\times d}$ is an invertible $d\times d$ matrix.
@@
$b\in C^{2}({\mathds{R}^{d}},{\mathds{R}^{d}})$ and there exist constants $\theta_{0}>0$, $K\geq 0$, $\theta_{1}\geq 0$, $\theta_{2}\geq 0$ such that for all $x,y,v,v_{1},v_{2}\in{\mathds{R}^{d}}$,
(1.3) $\langle x-y,b(x)-b(y)\rangle$ $\leq-\theta_{0}|x-y|^{2}+K,$
@@
(1.5) $|\nabla_{v_{2}}\nabla_{v_{1}}b(x)|$ $\leq\theta_{2}|v_{1}||v_{2}|.$
It is well known that under (1.4) the both SDEs (1.1) and (1.2) have unique non-explosive (strong) solutions, which we denote by $X^{x}_{t}$ and $Y^{y}_{t}$ respectively. From the classical Lyapunov function criterion, see e.g. [11], we know that under (H1) and (H2) the solutions to the SDEs (1.1) and (1.2) are ergodic. Denote by $\mu_{\alpha}$ and $\mu_{2}$ the respective ergodic measures of $X^{x}_{t}$ and $Y^{y}_{t}$.
Continuing the work in Continuing the work in [5], we establish an optimal bound between $X_{t}^{x}$ and $Y_{t}^{y}$ in total variation distance and consequently an optimal bound between $\mu_{\alpha}$ and $\mu_{2}$ as $t\rightarrow\infty$. Combining this bound with a recent interpolation result from [8], we can extend these bounds to the Wasserstein-$p$ distance with $0<p<1$.
1.1. Notation
Before giving our main results, let us first introduce the notation that will be used in sequel.
We denote by $C({\mathds{R}^{d}},\real^{m})$ the set of all continuous functions from $\mathds{R}^{d}$} to m and by $C_{b}({\mathds{R}^{d}},\real)$ the set of all bounded continuous functions from $\mathds{R}^{d}$} to . Given $k\in\mathds{N}$, denote by $C^{k}({\mathds{R}^{d}},\real)$ the set of all continuous functions from $\mathds{R}^{d}$} to with continuous 11st,…,$k$-th order derivatives, and by $C^{k}_{b}({\mathds{R}^{d}},\real)$ the set of all bounded continuous functions from $\mathds{R}^{d}$} to with bounded continuous 11st,…,$k$-th order derivatives. Let $f:{\mathds{R}^{d}}\to\real$ be sufficiently smooth, the directional derivative in direction $v\in{\mathds{R}^{d}}$ is defined as
@@ 1.1. Notation
and $\nabla_{v_{2}}\nabla_{v_{1}}f(x):=\nabla_{v_{2}}(\nabla_{v_{1}}f)(x)=\nabla^{2}f(x){v_{1}}\cdot{v_{2}}$ for $v_{1},v_{2}\in{\mathds{R}^{d}}$. The vector and matrix norms of $\nabla f$ and $\nabla^{2}f$ are given by
$|\nabla f(x)|=\sup_{v\in{\mathds{R}^{d}},|v|=1}\left|\nabla_{v}f(x)\right|$
and
$\|\nabla^{2}f(x)\|_{\mathrm{op}}=\sup_{v_{1},v_{2}\in{\mathds{R}^{d}},|v_{1}|=|v_{2}|=1}\left|\nabla_{v_{2}}\nabla_{v_{1}}f(x)\right|=\sup_{v_{1},v_{2}\in{\mathds{R}^{d}},|v_{1}|=|v_{2}|=1}\left\langle\nabla^{2}f(x),v_{1}v_{2}^{\top}\right\rangle_{\rm HS}.$
Moreover, we define
$\|\nabla f\|_{\infty}:=\sup_{x\in{\mathds{R}^{d}}}\|\nabla f(x)\|_{\mathrm{op}},\ \ \ \|\nabla^{2}f\|_{\mathrm{op},\infty}:=\sup_{x\in{\mathds{R}^{d}}}\|\nabla^{2}f(x)\|_{\mathrm{op}}.$
This notation extends naturally to vector-valued functions $f=(f_{1},f_{2},\cdots,f_{d})^{\top}:{\mathds{R}^{d}}\to{\mathds{R}^{d}}$:
@@
$\nabla^{2}f(x)v_{1}v_{2}$ $=\nabla_{v_{2}}\nabla_{v_{1}}f(x)=(\nabla_{v_{2}}\nabla_{v_{1}}f_{1}(x),\dots,\nabla_{v_{2}}\nabla_{v_{1}}f_{d}(x))^{\top}.$
For $x\in{\mathds{R}^{d}}$, we see $\nabla f(x)\in{\mathds{R}^{d}}$ and $\nabla^{2}f(x)\in\real^{d\times d}$. The Hilbert–Schmidt inner product of two matrices $A,B\in\real^{d\times d}$ is $\left\langle A,B\right\rangle_{\mathrm{HS}}=\sum_{i,j=1}^{d}A_{ij}B_{ij}$, and the Hilbert-Schmidt norm is $\|A\|_{\mathrm{HS}}=\sqrt{\sum_{i,j=1}^{d}A_{ij}^{2}}$.
Recall that $X_{t}^{x}$ and $Y_{t}^{y}$ are the solutions to SDEs (1.1) and (1.2) respectively. The operator semigroup induced by the Markov process $(X_{t}^{x})_{t\geq 0}$ is
@@
Write $\mathscr{C}_{b}(1)=\{h\in C_{b}({\mathds{R}^{d}},\real)\>;\>\|h\|_{\infty}\leq 1\}$ for the unit ball in $C_{b}({\mathds{R}^{d}},\real)$. The total variation distance of two probability measures (on $\mathds{R}^{d}$}) $\mu$ and $\nu$ is given by
(as usual we write $\mu(h):=\int h\,\mathrm{d}\mu$ etc.). It is clear that $\|\mu-\nu\|_{\mathrm{TV}}\leq 2$ for all probability measures $\mu$ and $\nu$. For $p\in(0,1]$, the Wasserstein-$p$ distance for two probability measures $\mu,\nu$ is given by
$\|\mu-\nu\|_{W_{p}}=\sup_{|h(x)-h(y)|\leq|x-y|^{p}}|\mu(h)-\nu(h)|.$
Finally, $\mathrm{Lip}(1)$ is the family of Lipschitz functions $f:{\mathds{R}^{d}}\to\real$ with Lipschitz constant $1$ and norm $\|f\|_{\mathrm{Lip}}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|}\leq 1$.
@@ 1.2. Main results
Throughout this paper, $C,C_{1},C_{2}$ denote positive constants which may depend on $\theta_{0},\theta_{1},\theta_{2},K,\|\sigma\|_{\mathrm{op}},\|\sigma^{-1}\|_{\mathrm{op}}$, but they are always independent of $d$ and $\alpha$; their values may change from line to line. Recall that $\mu_{\alpha}$ and $\mu_{2}$ are respective ergodic measures of $X^{x}_{t}$ and $Y^{y}_{t}$ respectively.
@@
$\qquad\leq C_{1}\left\{t^{-1/2}\vee 1\right\}\mathrm{e}^{-C_{2}t}|x-y|+\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
@@
(1.6) $\|\mu_{\alpha}-\mu_{2}\|_{\mathrm{TV}}\leq\frac{Cd\log(1+d)}{\alpha-1}(2-\alpha).$
We shall show that the bound in (1.6) is optimal with respect to $\alpha$ by a one-dimensional Ornstein–Uhlenbeck system, cf. Proposition 5.1 below.
Applying a recent interpolation result established in [8], we obtain the following corollary.
Corollary 1.2.
Assume (H1) and (H2). For any $p\in(0,1)$ and $\alpha\in(1,2)$,
$\|\mu_{\alpha}-\mu_{2}\|_{W_{p}}\leq\frac{Cd^{(p+3)/2}\log(1+d)}{\alpha-1}(2-\alpha).$
2. A key gradient estimate for Theorem 1.1
This section is devoted to proving the following gradient estimate by an integration by parts in Malliavin calculus, which is crucial for the proof of our main result.
@@
Assume (1.4) and (1.5). For any $t\in(0,1]$, $\alpha\in(0,2)$ and $f\in C^{2}_{b}({\mathds{R}^{d}},\real)$,
$\|P_{t}\nabla^{2}f\|_{\mathrm{op},\infty}\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\|\nabla f\|_{\infty}.$
@@ 2.1. A random time-change method
Let $\gamma^{\epsilon}$ be the inverse function of $\ell^{\epsilon}$, i.e., $\ell^{\epsilon}\circ\gamma^{\epsilon}_{t}=t$ for all $t\geq\ell_{0}^{\epsilon}$ and $\gamma^{\epsilon}\circ\ell^{\epsilon}_{t}=t$ for all $t\geq 0$. By construction, $t\mapsto\gamma_{t}^{\epsilon}$ is absolutely continuous on $\ell_{0}^{\epsilon},\infty)$. Define
@@ 2.2. Integration by parts via Malliavin Calculus
[…]
@@
Lemma $2.$
Assume (1.4) and (1.5), and fix $\ell\in\mathds{S}$ and $\epsilon\in(0,1)$. For any $v\in\real^{d}$, $t>0$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell^{\epsilon}}_{t}\right)\right]\right|\leq C|v|\left(\frac{1}{\sqrt{\ell^{\epsilon}_{t}-\ell^{\epsilon}_{0}}}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell^{\epsilon}}_{t})\right|^{2}\right]}.$
@@
For any fixed $t>\ell^{\epsilon}_{0}$ and $v\in\real^{d}$, let
(2.7) $u_{0}(s):=\frac{1}{t-\ell^{\epsilon}_{0}}\,\sigma^{-1}J_{s,t}^{-1}v.$
@@
$D_{u_{0}}Z^{x;\ell^{\epsilon}}_{t}=v.$
Combining this with the duality (2.5) between the Malliavin derivative and the divergence operator gives us, for any $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$,
$\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]$ $=\mathds{E}^{\mathds{W}}\left[\big\langle\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),v\big\rangle\right]=\mathds{E}^{\mathds{W}}\left[\big\langle\nabla f\left(Z^{x;\ell^{\epsilon}}_{t}\right),D_{u_{0}}Z^{x;\ell^{\epsilon}}_{t}\big\rangle\right]$
$=\mathds{E}^{\mathds{W}}\left[D_{u_{0}}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]=\mathds{E}^{\mathds{W}}\left[f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\delta\left(u_{0}\right)\right].$
By the Cauchy–Schwarz inequality and [13, Proposition 1.3.1], we obtain
(2.8) $\begin{split}\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|&\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\delta(u_{0})^{2}\right]}\\ &\leq\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right|^{2}\right]}\sqrt{\mathds{E}^{\mathds{W}}\left[\|u_{0}\|_{H}^{2}+\|Du_{0}\|_{H\otimes H}^{2}\right]},\end{split}$
@@
(2.9) $\begin{split}\|u_{0}\|_{H}^{2}&=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\exp\left(-\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)v\right|^{2}\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}|v|^{2}}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\exp\left(2\theta_{1}\int_{s}^{t}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}s\\ &\leq\frac{\|\sigma^{-1}\|_{\mathrm{op}}^{2}|v|^{2}}{t-\ell^{\epsilon}_{0}}\,\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}.\end{split}$
Similarly, using (1.4) and (2.6) gives for any $u\in H$
(2.10) $\left|D_{u}Z^{x;\ell^{\epsilon}}_{t}\right|\leq\|\sigma\|_{\mathrm{op}}\|u\|_{H}\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{t-\ell^{\epsilon}_{0}}.$
@@
$\begin{split}D_{u}J_{s,t}^{-1}&=\int_{0}^{1}\exp\left(-\tau\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\left(-\int_{s}^{t}\nabla^{2}b\left(Z^{x;\ell^{\epsilon}}_{r}\right)D_{u}Z^{x;\ell^{\epsilon}}_{r}\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\mbox{}\times\exp\left(-(1-\tau)\int_{s}^{t}\nabla b\left(Z^{x;\ell^{\epsilon}}_{r}\right)\dot{\gamma}^{\epsilon}_{r}\,\mathrm{d}r\right)\mathrm{d}\tau,\end{split}$